How do I find the vertices and foci of a parabola? I just need the formula; I can't find it in my textbook.
Answer by lwsshak3(11628)

How do I find the vertices and foci of a parabola? I just need the formula; I can't find it in my textbook.
***
There are 4 basic equations for a parabola:
(x-h)^2=4p(y-k) (parabola opens upward)
(x-h)^2=-4p(y-k) (parabola opens downward)
(y-k)^2=4p(x-h) (parabola opens rightward)
(y-k)^2=-4p(x-h) (parabola opens leftward)
In every case, (h,k)=(x,y) coordinates of the vertex
Focus=p-units from the vertex on the axis of symmetry
p=distance from vertex to focus and directrix on the axis of symmetry.